OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL: 1. Brianna plotted the two points (20, 75) and (45, 150) on a graph. 1. Write the equation of the line in slope-intercept form. 2. What is the x-coordinate of the x-intercept of the line that contains these two points?

OpenStudy (anonymous):

@johnweldon1993 do you think you could help me

OpenStudy (johnweldon1993):

Well first we need to find the slope between these 2 points \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

I got -3 as the slope

OpenStudy (johnweldon1993):

Why -3? \[\large \frac{150 - 75}{45 - 20}\]

OpenStudy (anonymous):

oh, I did it backwards

OpenStudy (anonymous):

it's positive 3

OpenStudy (johnweldon1993):

Correct...now to put the points into slope intercept form...we just use the point slope formula \[\large y - y_0 = m(x - x_0)\] Pick which point you want to use...

OpenStudy (anonymous):

\[y - 75 = 3(x - 20)\]\[y - 75 = 3x - 60\]\[y = 3x + 15\]

OpenStudy (johnweldon1993):

Correct! And finally what is the x-coordinate of the intercept?

OpenStudy (anonymous):

is it 15

OpenStudy (johnweldon1993):

Hmm are you sure? The X intercept occurs when y = 0

OpenStudy (johnweldon1993):

15 would be the 'y' intercept

OpenStudy (anonymous):

\[0 = 3x + 15 \]\[-15 = 3x \]\[-5 = x \] I did this equation... and I got -5 idk if this is correct or not

OpenStudy (johnweldon1993):

That is exactly correct

OpenStudy (anonymous):

Great, thanks for your help :)